OpenStudy (anonymous):

calculate the product and express the answer in the form a+bi {cos(pi/12) + isin(pi/12)}^6

OpenStudy (anonymous):

multiply the angle by 6 and then convert

OpenStudy (anonymous):

you get \[\cos(\frac{\pi}{2})+i\sin(\frac{\pi}{2})\]

OpenStudy (anonymous):

then since \(\cos(\frac{\pi}{2})=0\) and \(\sin(\frac{\pi}{2})=1\) you are left with \(i\)

OpenStudy (anonymous):

or if you want to be silly you can write \(0+i\) hope it is clear that this is actually rather easy, since all it requires is multiplying the angle by 6

OpenStudy (anonymous):

thanks